# UniMS-RAG: A Unified Multi-source Retrieval-Augmented Generation for Personalized Dialogue Systems

### v1

@@ Abstract.
Large Language Models (LLMs) have shown exceptional capabilities in many natural language understanding and generation tasks. However, the personalization issue still remains a much-coveted property, especially when it comes to the multiple sources involved in the dialogue system. To better plan and incorporate the use of multiple sources in generating personalized response, we firstly decompose it into three sub-tasks: Knowledge Source Selection, Knowledge Retrieval, and Response Generation. We then propose a novel Unified Multi-Source Retrieval-Augmented Generation system (UniMS-RAG) Specifically, we unify these three sub-tasks with different formulations into the same sequence-to-sequence paradigm during the training, to adaptively retrieve evidences and evaluate the relevance on-demand using special tokens, called acting tokens and evaluation tokens. Enabling language models to generate acting tokens facilitates interaction with various knowledge sources, allowing them to adapt their behavior to diverse task requirements. Meanwhile, evaluation tokens gauge the relevance score between the dialogue context and the retrieved evidence. In addition, we carefully design a self-refinement mechanism to iteratively refine the generated response considering 1) the consistency scores between the generated response and retrieved evidence; and 2) the relevance scores. Experiments on two personalized datasets (DuLeMon and KBP) show that UniMS-RAG achieves state-of-the-art performance on the knowledge source selection and response generation task with itself as a retriever in a unified manner. Extensive analyses and discussions are provided for shedding some new perspectives for personalized dialogue systems.
@@
Dialogue System, Retrieval-Augmented Generation, Large Language Models, Personalized Response Generation
This article substantially extends our previous work published as a conference paper in EMNLP 2023 (Wang et al., 2023a). The conference version proposed KBP dataset and SAFARI framework to utilize LLMs as source planner and reader. Compared with the previous version, this manuscript propose UniMS-RAG: a unified multi-source retrieval-augmented generation framework, which 1) introduce additional sub-task: relevance score predication with carefully designed evidence attention mask during the training, enabling itself to serve as planner, retriever and reader in a unified manner. We also explore the effects of different relevance signals, including DPR and LLMs prompting; and 2) contains self-refinement mechanism to iteratively refine the generated response during the inference, considering its relationship between the retrieved evidence (consistency scores) and also the relationship between the dialogue context and retrieved evidence (relevance scores). In addition, we conduct experiments on more personalized datasets to validate the robustness and effectiveness of UniMS-RAG. The proposed framework is further evaluated through extensive experiments in-depth analyses, including automatic evaluation and human evaluation. To conclude, our proposed UniMS-RAG not only substantially modify the training and inference of SAFARI, but also can be more extensible and preferable in personalized dialogue system for better performance and flexibility.
This paper was partially supported by grants from the RGC General Research Funding Scheme (GRF) 14222922 (CUHK 2151185). Work done when Hongru Wang is remotely visiting The University of Edinburgh.
Author Address: H. Wang, Z. Wang, K.-F. Wong (corresponding author), The Chinese University of Hong Kong, Hong Kong, 999077; emails: {hrwang, zzwang, kfwong}@se.cuhk.edu.hk; W. Huang, J.Z. Pan (corresponding author), The University of Edinburgh, United Kingdom; emails: {w.huang, j.z.pan}@ed.ac.uk; Y. Deng, National University of Singapore, Singapore; email: ydeng@nus.edu.sg; R. Wang, Harbin Institute of Technology, Shenzhen, China; email: ruiwangnlp@outlook.com; Y. Wang, F. Mi, Huawei Noah Ark Lab, China, emails: {wangyufei44, mifei2}@huawei.com
copyright: acmcopyrightjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYprice: 15.00isbn: 978-1-4503-XXXX-X/18/06ccs: Information systems Personalizationccs: Computing methodologies Discourse, dialogue and pragmaticsccs: Computing methodologies Planning for deterministic actionsccs: Computing methodologies Natural language generationccs: Information systems Language models
1. Introduction
@@
The emergence of large language models (LLMs) has revolutionized the field of natural language processing, including many downstream understanding and generation tasks (Bang et al., 2023). While these models have undeniably advanced the state of the art in various applications, they also introduce new challenges, particularly the factual error (Mallen et al., 2023; Xue et al., 2023) and personalization issues (Salemi et al., 2023; Wang et al., 2023d) in the realm of dialogue systems. To alleviate this, Retrieval-Augmented Generation (RAG) methods are usually adopted to retrieve relevant passages, aiming to enrich the semantic information of the dialogue context, resulting in up-to-date, factual and personalized responses. Furthermore, the latest Self-RAG (Asai et al., 2023b) considers the cases that require external knowledge and the cases that do not at the same time, achieving a better trade-off between effectiveness and efficiency. In a world inundated with vast amounts of information, retrieval-augmented dialogue systems play a crucial role in ensuring that automated conversations are not only linguistically proficient but also factually reliable and contextually aware.
(a) An example that the response requires both user persona and system persona information at DuLeMon.
(b) An example that the response requires both user persona and corresponding documents at KBP. In this way, the retriever need to firstly retrieve related persona (1st step), and then retrieve the document according to dialogue context and retrieved persona information (2nd step).
Figure 1. Two typical examples of multi-source personalized knowledge-grounded dialogues: a): An example from DuLeMon (Xu et al., 2022a); and b) An example from KBP (Wang et al., 2023a). We use same color to indicate the response and corresponding grounded knowledge. We skip the dialogue context for simplicity.
@@
There are two limitations. On the one hand, most of the existing knowledge-grounded dialogue systems either focus on a single source of knowledge or indiscriminately incorporate all sources of knowledge. In detail, previous methods usually interact with different external sources of knowledge to engage the user and provide human-like conversational experience, such as system persona source to maintain consistency in responses (Xu et al., 2022a), user memory or profile source for personalized responses (Madotto et al., 2019; Salemi et al., 2023), Wikipedia (Dinan et al., 2019a) or the internet (Komeili et al., 2022) source to ensure the provision of up-to-date information. Yet, they focus on individual sources or tasks in isolation, overlooking the complexity and the potential existence of multiple sources of knowledge in real-world scenarios, as shown in Figure 1. In this way, the dialogue system needs to decide which source to use (or not use) dynamically, in order to provide more personalized and helpful responses. Furthermore, the inter-dependency relationship between these different sources of knowledge needs to be considered to guide better response generation (Wang et al., 2023a). For example, as shown in Figure 0(b), the dialogue system needs to locate the specific persona config first according to the dialogue context, and then find relevant dependent documents to answer the user inquiry. On the other hand, previous work either independently train the retriever and reader (Wang et al., 2023a; Ram et al., 2023), resulting in sub-optimal performance and distribution shift between retriever and reader; or design complex architecture to optimize them at the same time (Guu et al., 2020a; Rubin and Berant, 2023), which is infeasible and inefficient at the era of large language models due to unaffordable computing cost.
@@
Then we design a novel framework, Unified Multi-Source Retrieval-Augmented Dialogue System (UniMS-RAG), that unifies three tasks in the above using the same large language models in a sequence-to-sequence (Seq2Seq) manner. In specific, motivated by recent work of assigning different tokens with different roles (Wang et al., 2023a; Asai et al., 2023b), we carefully introduce two types of tokens: 1) acting tokens to decide the next action (e.g., which source to use); and 2) and evaluation tokens to evaluate the relevance score between dialogue context and retrieved evidence (e.g., the similarity score). Thus we can reformulate the above three tasks as token prediction tasks by generating acting tokens (planner), evaluation tokens (retriever) or original tokens in the vocabulary (reader) during the training. To further enhance the quality of generated responses, we incorporate a self-refinement process during the inference stage. This involves reassessing the generated responses by leveraging feedback from evaluation tokens and ensuring consistency between the provided evidence and the responses. To sum up, the contributions are summarized as follows:
  • We formally propose a multi-source personalized knowledge-grounded dialogue tasks, consisting of three different sub-tasks: knowledge source selection, knowledge retrieval and final response generation.
@@
  • We investigate the different strategies to get the soft label of evaluation tokens during the training stage, including prompting LLMs or using an independent fine-tuned retriever. Furthermore, we propose a self-refinement mechanism to re-generate the response using updated evidence according to its relevance with the dialogue context and previously generated response.
@@
  • Experimental results on two PerDS benchmark datasets show that UniMS-RAG achieves state-of-the-art performance on the knowledge selection and response generation tasks with itself serving as a retriever, resulting in more personalized and factual responses. Extensive analyses provide some new insights into the future of multi-source retrieval-augmented generation tasks.
2. Related Work
2.1. Personalized Dialogue System
@@
To build a personalized dialogue agent, Zhang et al. (2018) firstly investigated this task with a new dataset Persona-Chat, where a pre-defined persona set is a form of multiple sentences of textual description. Lots of works follow this setting and have taken mutual persona perception (Liu et al., 2020; Kim et al., 2020; Xu et al., 2022c), persona-sparse scenario (Song et al., 2021; Welch et al., 2022; Song et al., 2023), long-term persona memory (Xu et al., 2022a), persona extending (Liu et al., 2022) and persona order bias (Chen et al., 2023) into consideration. Although some of them complement the insufficient semantics in short persona descriptions by further utilizing an external commonsense knowledge base to extend existing persona sets (Majumder et al., 2020; Liu et al., 2022), they still fall into the conventional framework coupling the knowledge selection with the response generation (Wu et al., 2022b), rendering it infeasible to handle various sources of knowledge. There have also been works showing that the combination of different knowledge sources such as persona descriptions and Wikipedia can further improve the overall performance (Jang et al., 2022a; Wu et al., 2021, 2022a). However, they still fail to capture possible dependency between knowledge sources. In their framework, knowledge is not used as the role to assist persona-consistent response generation, but as an additional resource to generate a more informative response (Dinan et al., 2019b; Xue et al., 2023) or select a suitable persona (Jang et al., 2022a; Fu et al., 2022). Furthermore, most existing works overlook the possibilities that the response does not require the involvement of persona descriptions by simply concatenating all personas with the dialogue context to generate the final response (Majumder et al., 2020; Song et al., 2023; Xu et al., 2022c).
2.2. Knowledge-grounded Dialogue System
@@
How to interact with different external sources plays a key role in dialogue systems to retrieve corresponding knowledge, resulting in more helpful, personalized, trustworthy responses (Huang et al., 2020; Wang et al., 2023c). Specifically, most of previous methods rely on different external sources of knowledge to engage the user and improve the conversational experience, including but not limited to system persona to maintain consistency in responses (Wang et al., 2023a), user memory or profile for personalized interactions (Xu et al., 2022b; Zhang et al., 2019), and access to sources like Wikipedia or the internet to ensure the provision of up-to-date information (Meng et al., 2020; Nakano et al., 2022). However, most of them focus on individual sources or tasks in isolation, overlooking the complexity and the potential existence of multiple sources in practice. There is a latest work named TPE which regards different knowledge sources as conceptual tools and proposes a multi-persona collaboration framework to model the decision-making process of the call order for multiple knowledge sources (Wang et al., 2023e). We differ in exploring the capability of LLMs to be planner, retriever, reader in a unified manner.
2.3. Retrieval-augmented Generation
@@
Retrieval-augmented Generation (RAG) has been considered as an effective method to overcome several limitations of LLMs, such as hallucinations (Shuster et al., 2021; Ji et al., 2023), factuality (Wang et al., 2023b), long-term memory (Xu et al., 2022a) and etc (Gao et al., 2024). Usually, an external retriever is first used to retrieve relevant textual knowledge from one specific knowledge source (e.g., Wikipedia), then the reader takes the relevant textual knowledge as external context for generating knowledge-grounded response (Lewis et al., 2020). Most of previous works try to optimize the retriever and reader independently (Gao et al., 2024). During the initial phases, people use sparse retriever, such as BM25 (Robertson et al., 2009) and TF-IDF (Wu et al., 2008), to make relevance decisions and retrieve corresponding evidence. However, sparse approaches fall short in extracting the semantic features inherent in text content (Guo et al., 2022). To overcome this issue, researchers have proposed language model-based dense retrieval methods by encoding documents and queries as dense vectors, which effectively represent the semantic features of text content (Karpukhin et al., 2020; Li et al., 2023; Bruch et al., 2023). As the first dense retriever, DPR (Karpukhin et al., 2020) uses two pre-trained language models to encode documents and queries separately, allowing for a more nuanced understanding of the content. In addition, Glass et al. (2022) propose a retrieve and re-rank framework for leveraging both advantages of sparse retrieval and dense retrieval. There are also several attempts which optimize the retriever and reader simultaneously (Asai et al., 2023a). For example, Guu et al. (2020b) augment language model pretraining with a latent knowledge retriever, allowing the model to retrieve and attend over documents used during pre-training, fine-tuning and inference. However, the complex architectures and interactions make them infeasible and inefficient in the era of LLMs.
More recently, there are a handful of works exploring the performance of LLMs as retriever (Zhu et al., 2023; Sun et al., 2023a; Shen et al., 2023; Ma et al., 2023). In detail, Shen et al. (2023) firstly prove that LLMs can be used as strong zero-shot retriever on several benchmark datasets, while Ma et al. (2023) propose Listwise Reranker with a Large Language Model (LRL), which achieves strong reranking effectiveness without using any task-specific training data. More recently, Sun et al. (2023a) investigate generative LLMs such as ChatGPT (Ouyang et al., 2022) and GPT-4  [^1] https://openai.com/research/gpt-4 for relevance ranking, and they surprisingly find that properly instructed LLMs can output competitive, even superior results to state-of-the-art supervised methods on popular information retrieval benchmarks. Additionally, they delve into the potential for distilling the ranking capabilities of ChatGPT into small specialized models, in order to improve efficiency in real-world applications. Distinguishing from previous works, we finetune LLM itself to learn a joint distribution of dialogue and evidence using similarity feedbacks from current most powerful LLMs, such as ChatGPT and GPT-4.
3. Problem Definition
@@
To consider the responses which require external knowledge and those which do not in practice, we provide a unified definition to unify the retrieval-augmented dialogue system and non-retrieval dialogue system, following Wang et al. (Wang et al., 2023a). Let $\boldsymbol{C}_{t}=\{u_{1},s_{1},u_{2},s_{2},...,u_{t}\}$ denotes the dialogue context at the current conversation turn $t$, and there are different knowledge sources $\boldsymbol{K}=\{K_{1},K_{2},...,K_{i}\}$, where $K_{i}=\{k_{i}^{1},k_{i}^{2},...,k_{i}^{j}\}$ indicates the $i_{th}$ source’s name of $\boldsymbol{K}$. $k_{i}^{j}$ denotes the $j_{th}$ knowledge in natural language from $K_{i}$. The goal of retrieval-augmented dialogue system [^2] In this context, a non-retrieval dialogue system is viewed as a specific type within retrieval-augmented dialogue systems, distinguished by the absence of a knowledge source, denoted as NULL. is to select suitable knowledge sources, and then generate the helpful, informative and personalized response, depending on which source is chosen (Wang et al., 2023c). Thus, the problem of PerDS can be decomposed into the following three tasks:

### v2

@@ Abstract
Large Language Models (LLMs) have shown exceptional capabilities in many natural language understanding and generation tasks. However, the personalization issue still remains a much-coveted property, especially when it comes to the multiple sources involved in the dialogue system. To better plan and incorporate the use of multiple sources in generating personalized response, we firstly decompose it into three sub-tasks: Knowledge Source Selection, Knowledge Retrieval, and Response Generation. We then propose a novel Unified Multi-Source Retrieval-Augmented Generation system (UniMS-RAG) Specifically, we unify these three sub-tasks with different formulations into the same sequence-to-sequence paradigm during the training, to adaptively retrieve evidences and evaluate the relevance on-demand using special tokens, called acting tokens and evaluation tokens. Enabling language models to generate acting tokens facilitates interaction with various knowledge sources, allowing them to adapt their behavior to diverse task requirements. Meanwhile, evaluation tokens gauge the relevance score between the dialogue context and the retrieved evidence. In addition, we carefully design a self-refinement mechanism to iteratively refine the generated response considering 1) the consistency scores between the generated response and retrieved evidence; and 2) the relevance scores. Experiments on two personalized datasets (DuLeMon and KBP) show that UniMS-RAG achieves better performance than previous strong baselines on the knowledge source selection and response generation task with itself as a retriever in a unified manner, and achieves new state-of-the-art when using more advanced external retriever. Extensive analyses and discussions are provided for shedding some new perspectives for personalized dialogue systems.
Index Terms:
@@
Open-domain Dialogue System, Large Language Models, Retrieval-Augmented Generation
1 Introduction
@@
Tthe emergence of large language models (LLMs) has revolutionized the field of natural language processing, including many downstream understanding and generation tasks [1]. While these models have undeniably advanced the state of the art in various applications, they also introduce new challenges, particularly the factual error [2, 3] and personalization issues [4, 5] in the realm of dialogue systems. To alleviate this, Retrieval-Augmented Generation (RAG) methods are usually adopted to retrieve relevant passages, aiming to enrich the semantic information of the dialogue context, resulting in up-to-date, factual and personalized responses. Furthermore, the latest Self-RAG [6] considers the cases that require external knowledge and the cases that do not at the same time, achieving a better trade-off between effectiveness and efficiency. In a world inundated with vast amounts of information, retrieval-augmented dialogue systems play a crucial role in ensuring that automated conversations are not only linguistically proficient but also factually reliable and contextually aware.
@@
There are two limitations. On the one hand, most of the existing knowledge-grounded dialogue systems either focus on a single source of knowledge or indiscriminately incorporate all sources of knowledge. In detail, previous methods usually interact with different external sources of knowledge to engage the user and provide human-like conversational experience, such as system persona source to maintain consistency in responses [7], user memory or profile source for personalized responses [8, 4], Wikipedia [9] or the internet [10] source to ensure the provision of up-to-date information. Yet, they focus on individual sources or tasks in isolation, overlooking the complexity and the potential existence of multiple sources of knowledge in real-world scenarios, as shown in Figure 1. In this way, the dialogue system needs to decide which source to use (or not use) dynamically, in order to provide more personalized and helpful responses. Furthermore, the inter-dependency relationship between these different sources of knowledge needs to be considered to guide better response generation [11]. For example, as shown in bottom part of Figure 1, the dialogue system needs to locate the specific persona config first according to the dialogue context, and then find relevant dependent documents to answer the user inquiry. On the other hand, previous work either independently train the retriever and reader [11, 12], resulting in sub-optimal performance and distribution shift between retriever and reader; or design complex architecture to optimize them at the same time [13, 14], which is infeasible and inefficient at the era of large language models due to unaffordable computing cost.
@@
Then we design a novel framework, Unified Multi-Source Retrieval-Augmented Dialogue System (UniMS-RAG), that unifies three tasks in the above using the same large language models in a sequence-to-sequence (Seq2Seq) manner. In specific, motivated by recent work of assigning different tokens with different roles [11, 6], we carefully introduce two types of tokens: 1) acting tokens to decide the next action (e.g., which source to use), aiming to call different source on-demand instead of incorporating all of them; and 2) and evaluation tokens to evaluate the relevance score between dialogue context and retrieved evidence (e.g., the similarity score), in order to force the model attention on more relevant evidence while overlooking noisy ones. Thus we can reformulate the above three tasks as token prediction tasks by generating acting tokens (planner), evaluation tokens (retriever) or original tokens in the vocabulary (reader) during the training. We randomly shuffle the order of retrieved evidences to prevent model learning a shortcut by attentions on evidences appears in specific positions. To further enhance the quality of generated responses, we incorporate a self-refinement process during the inference stage. This involves reassessing the generated responses by leveraging feedback from evaluation tokens and ensuring consistency between the provided evidence and the responses. To sum up, the contributions are summarized as follows:
  • We formally propose a multi-source personalized knowledge-grounded dialogue tasks, consisting of three different sub-tasks: knowledge source selection, knowledge retrieval and final response generation. Our analysis provides detailed insights into the limitations and proficiencies of current LLMs across these sub-tasks.
@@
  • We investigate the different strategies to get the soft label of evaluation tokens during the training stage, including prompting LLMs or using an independent fine-tuned retriever (a.k.a, classification-based and prompting-based methods). Furthermore, we propose a self-refinement mechanism to re-generate the response using updated evidence according to its relevance with the dialogue context and previously generated response.
@@
  • Experimental results on two PerDS benchmark datasets show that UniMS-RAG outperforms previous strong baselines, and achieves state-of-the-art performance with external more effective retriever, resulting in more personalized and factual responses. Extensive analyses provide some new insights into the future of multi-source retrieval-augmented generation tasks.
Figure 1: Two typical examples of multi-source personalized knowledge-grounded dialogues: upper): An example from DuLeMon [7]; and bottom): An example from KBP [11]. We use same color to indicate the response and corresponding grounded knowledge. We skip the dialogue context for simplicity.
2 Related Work
2.1 Personalized Dialogue System
@@
To build a personalized dialogue agent, Zhang et al. [15] firstly investigated this task with a new dataset Persona-Chat, where a pre-defined persona set is a form of multiple sentences of textual description. Lots of works follow this setting and have taken mutual persona perception [16, 17], persona-sparse scenario [18, 19], long-term persona memory [7], persona extending [20] and persona order bias [21] into consideration. Although some of them complement the insufficient semantics in short persona descriptions by further utilizing an external commonsense knowledge base to extend existing persona sets [22, 20], they still fall into the conventional framework coupling the knowledge selection with the response generation [23], rendering it infeasible to handle various sources of knowledge. There have also been works showing that the combination of different knowledge sources such as persona descriptions and Wikipedia can further improve the overall performance [24, 25, 26]. However, they still fail to capture possible dependency between knowledge sources. In their framework, knowledge is not used as the role to assist persona-consistent response generation, but as an additional resource to generate a more informative response [27, 3] or select a suitable persona [24, 28]. Furthermore, most existing works overlook the possibilities that the response does not require the involvement of persona descriptions by simply concatenating all personas with the dialogue context to generate the final response [22, 17].
2.2 Knowledge-grounded Dialogue System
@@
How to interact with different external sources plays a key role in dialogue systems to retrieve corresponding knowledge, resulting in more helpful, personalized, trustworthy responses [29, 30]. Specifically, most of previous methods rely on different external sources of knowledge to engage the user and improve the conversational experience, including but not limited to system persona to maintain consistency in responses [11], user memory or profile for personalized interactions [31, 32], and access to sources like Wikipedia or the internet to ensure the provision of up-to-date information [33, 34]. However, most of them focus on individual sources or tasks in isolation, overlooking the complexity and the potential existence of multiple sources in practice. There is a latest work named TPE which regards different knowledge sources as conceptual tools and proposes a multi-persona collaboration framework to model the decision-making process of the call order for multiple knowledge sources [35]. We differ in exploring the capability of LLMs to be planner, retriever, reader in a unified manner.
2.3 Retrieval-augmented Generation
@@
Retrieval-augmented Generation (RAG) has been considered as an effective method to overcome several limitations of LLMs, such as hallucinations [36], factuality [37] and long-term memory [7]. Usually, an external retriever is first used to retrieve relevant textual knowledge from one specific knowledge source (e.g., Wikipedia), then the reader takes the relevant textual knowledge as external context for generating knowledge-grounded response [38]. Most of previous works try to optimize the retriever and reader independently [39]. During the initial phases, people use sparse retriever, such as BM25 [40], to make relevance decisions and retrieve corresponding evidence. However, sparse approaches fall short in extracting the semantic features inherent in text content [41]. To overcome this issue, researchers have proposed language model-based dense retrieval methods by encoding documents and queries as dense vectors, which effectively represent the semantic features of text content [42, 43, 44]. For example, DPR [42] uses two pre-trained language models to encode documents and queries separately, allowing for a more nuanced understanding of the content. More recently, there are a handful of works exploring the performance of LLMs as retriever [45, 46, 47, 48]. In detail, Shen et al. [47] firstly prove that LLMs can be used as strong zero-shot retriever on several benchmark datasets, while Ma et al. [48] propose Listwise Reranker with a Large Language Model (LRL), which achieves strong reranking effectiveness without using any task-specific training data. Distinguishing from previous works, we finetune LLM itself to learn a joint distribution of dialogue and evidence using similarity feedbacks from current most powerful LLMs, such as ChatGPT and GPT-4.
3 Problem Definition
@@
To consider the responses which require external knowledge and those which do not in practice, we provide a unified definition to unify the retrieval-augmented dialogue system and non-retrieval dialogue system, following Wang et al. [11]. Let $\boldsymbol{C}_{t}=\{u_{1},s_{1},u_{2},s_{2},...,u_{t}\}$ denotes the dialogue context at the current conversation turn $t$, and there are different knowledge sources $\boldsymbol{K}=\{K_{1},K_{2},...,K_{i}\}$, where $K_{i}=\{k_{i}^{1},k_{i}^{2},...,k_{i}^{j}\}$ indicates the $i_{th}$ source’s name of $\boldsymbol{K}$. $k_{i}^{j}$ denotes the $j_{th}$ knowledge in natural language from $K_{i}$. The goal of retrieval-augmented dialogue system [^1] In this context, a non-retrieval dialogue system is viewed as a specific type within retrieval-augmented dialogue systems, distinguished by the absence of a knowledge source, denoted as NULL. is to select suitable knowledge sources, and then generate the helpful, informative and personalized response, depending on which source is chosen [30]. Thus, the problem of PerDS can be decomposed into the following three tasks: